(7,9)-(3,-7)-(-3,3) are the points 
equaliteral triangle
issoscles triangle
right tiangled
issoscle right triangle
 

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
let A=(7,9) B=(3,-7) and C=(-3,3)
AB = sqrt((3-7)^2 +(-7-9)^2) =sqrt(272)
similarly, BC=sqrt(136) and CA=sqrt(136)
 
as CA=BC, so it is an isosceles triangle.
now BC^2 + CA^2= 272=AB^2 this satisfies Pythagoras’ Theorem
So it is right angled.
Therefore the triangle is a right angled isosceles triangle. [Ans]
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)